OpenStudy (anonymous):

Use the matrix method to solve 2x + y = 9

OpenStudy (mathmale):

This is just a single equation, whereas the matrix method is used to solve SYSTEMS of linear equations. Are you certain that you've posted the entire problem?

OpenStudy (anonymous):

yes all they want to know y=

OpenStudy (mathmale):

I doubt that you'd need matrices to solve this equation for y, and also doubt that matrices apply to such a problem. 2x + y = 9 can be solved for y by subtracting 2x from both sides and simplifying the result. Please try this. If you have an example of "using the matrix method to solve 2x+y-9," please share it.

OpenStudy (anonymous):

ok then what is y=1

OpenStudy (unklerhaukus):

what are the two equations in the question,

OpenStudy (anonymous):

x + 4y = 8 2x + y = 9

OpenStudy (unklerhaukus):

\[x + 4y = 8\\ 2x + y = 9\] \[\left[ \begin{array}{cc}1&4\\ 2&1\end{array}\right]\left[\begin{array}{c}x\\y \end{array}\right] =\left[\begin{array}{c}8\\9 \end{array}\right]\] \[\left[\begin{array}{c}1&4\\2&1 \end{array}\left|\begin{array}{c}8\\9 \end{array}\right.\right]\]

OpenStudy (unklerhaukus):

so reduce the matrix, start by changing the second row (\(R_2\)), to become \(R_2-2R_1\) ,

OpenStudy (mathmale):

@J14212 : As I had suspected, you did not include all of the information that came with this math problem in your initial statement. Please, to save yourself and potential helpers time and effort, double-check that your problem statement is complete before you click on that Post button. x + 4y = 8 2x + y = 9 Uncle Rhaukus is using the "Gauss-Jordan Elimination Method" to solve this system of linear equations. That's a great method, but there are other methods as well.